Show your solution.




1. When a 12.0-V car battery powers a single 30.0 W headlight, how many electrons pass through it each second?




2. Two electrons are 10^5 m apart. One electron is fixed on its position. The other, originally at rest, is released. How fast is the free electron moving when it is 10^ βˆ’3 m from the fixed charge? Apply conservation of energy.




3. Two point charges, +400 uC and -2.00 uC are 0.55m apart. (a) what is the potential energy of the pair? (b) If the positive charge is fixed and the other is released from rest, how far from each other should the charges be if the negative charge is moving with a speed of 2.30 x 10^2 m/s? Let the mass of the moving charge be 8.00 x 10^βˆ’6 kg.

1
Expert's answer
2022-04-13T09:21:50-0400

1.

"U=\frac Aq=\frac{Pt}{ne},"

"n=\frac{Pt}{eU}=1.6\cdot10^{19}."

2.

"U=\frac{kq_1q_2}r=13~J."

3.

"U=\frac{mv^2}2+\frac{kq_1q_2}d,"

"d=56~cm."